Construct a one-dimension dynamical systems F(x) so that it has 2 three equilibrium points. Moreover, two of them are semi-stHowdy, any help with this here would be a blessing. Thank ya'll!
Answer #1

Solution: 2. Definition: An equilibrium solution is any constant (horizontal) function x(t) = c that is a solution to the differential equation.

Stable: The equilibrium solution x(t) = c is stable if all solutions with initial conditions x_0 `near' x = c approach c as t \to \infty.
Unstable: The equilibrium solution x(t) = c is unstable if all solutions with initial conditions x_0 `near' x= c do n aotpproach c as t \to \infty.
Semi-stable: The equilibrium solution x(t) = c is semi-stable if initial conditions x_0 on one side of c lead to solutions x(t) that approach c as t \to \infty , while initial conditions x_0 on the other side of c do not approach c.

Construction of one-dimension dynamical system that has three equilibrium points:

(a) Let

d.x lt

(i) Here equilibrium points are given by

\frac{dx}{dt}=x'=-(x-5)(x-8)^2=F(x)=0

\Rightarrow(x-5)(x-8)^2=0

\Rightarrow x=5, x=8,8

Therefore, r(t) 5 and (t)-8, 8 are three equilibrium solutions.

(ii) For, d.x lt is negative, so x(t) is decreasing.

For, dt is negative, so x(t) is decreasing.

For, dt is positive, so x(t) is increasing.

Thus x(t) = 5 is stable and x(t)= 8 is semi-stable.

(b) Let

d.x lt

Here x=4 is stable and x=6 and x=9 are semistable